OpenStudy (anonymous):

Eliminate the “z” variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 Answer 6x – 5y = –13 4x – 3y = –14 6x + 5y = 13 4x + y = –14

OpenStudy (anonymous):

Last two equations are 2x – y + 2z = –15 <---- Here coefficient of z is +2 and its opposite is -2 2x – 2y – z = 1 <----- If we multiply this equation by 2, then -z becomes -2z so then we get 2x – y + 2z = –15 4x – 4y – 2z = 2 Now on adding the two we get.. 6x - 5y = -13 so answer is the first choice 6x – 5y = –13